Session: Nov 2003 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Sun Nov 9 03:59:27 2003) Success :
Everything OK

Shanghai (Xinyong Huang at Sun Nov 9 03:59:53 2003) Success :

Urumqi (zhanghua at Sun Nov 9 04:41:24 2003) Success :
From 10:03 to 18:30 and from 21:03 to 23:35,the source is outside telescope limits.

Onsala (Michael Lindqvist at Sun Nov 9 08:59:20 2003) Success :
No known problems.

Jodrell_Bank (Alastair Gunn at Mon Nov 10 00:43:30 2003) Minor failures : RECO
Problems with tape at start. Schedule started at scan 9 at 04:55 UT. Schedule restarted at 19:25 and 20:05 due to tape problems. Apparently no loss of data. No other known problems.

Westerbork (tony foley at Mon Nov 10 10:06:51 2003) Minor failures : RECO
problems starting parallel Mk5 recordings on a new diskpack. Seems like it needs "erase=reset" with a blank pack in A. Parallel Mk5 started at 05:19UT. Second pack starts at 21:20UT.

Effelsberg (A. Kraus at Mon Nov 10 10:29:43 2003) Minor failures : RECO HIGH
Due to MK5 problems, not data was recorded between UT0400-0444. Few Tsys measurements (on 3C274) were not ok.

Noto (P. Cassaro at Mon Nov 10 16:26:26 2003) Minor failures : RECO
Data lost from 03:25 UT to the end, due to vacuum problems.